# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ 3. Phase 1: Development of psychological risks taxonomy
We divided our study into two phases (see Figure 1). In Phase 1, we developed a comprehensive psychological risk taxonomy associated with AI conversational agents, informed by survey responses and validated through workshop sessions in Phase 2. This section outlines our approach for structuring psychological risks, survey design, and survey results. We also presents the finalized psychological risk taxonomy, which includes potentially harmful AI behaviors, negative psychological impacts, and contextual factors affecting individuals. Finally, we provide case studies illustrating the interplay between AI behaviors and psychological impacts, highlighting how specific contexts mediate or exacerbate these effects.
@@ 3.1. Method
In the first phase, we conducted a survey study with 283 participants to gather a broad range of in-the-wild experiences with AI conversational agents, specifically focusing on experiences that led to negative psychological impacts. In this subsection, we start with describing the process for structuring psychological risks, followed by survey participant and design details. We end the subsection with qualitative analysis of the survey responses.

### v3

@@ 3. Phase 1: Development of psychological risks taxonomy
We divided our study into two phases (Figure 1). In Phase 1, we developed a comprehensive psychological risk taxonomy for AI conversational agents, informed by survey responses and validated through workshop sessions in Phase 2. This section focuses on Phase 1, including our methodology (Section 3.1), the finalized psychological risk taxonomy (Section 3.2), and case studies illustrating the interplay between behaviors, impacts, and contexts (Section 3.3). Both phases of the study were reviewed and approved by the institution’s ethics review board.
@@ 3.1. Method
We conducted a survey study with 283 participants to gather a broad range of in-the-wild experiences with AI conversational agents.